3. The number of goldfish in a tank is 22, and the volume of the tank is 56 cubic feet. What is the density of the tank?


0.34 goldfish per cubic foot

0.39 goldfish per cubic foot

1.41 goldfish per cubic foot

2.55 goldfish per cubic foot




4. The population of Tinyville is 16.2 Tinys per acre. Exactly 405 Tinys live in Tinyville. How many acres is Tinyville? Only enter a numerical value in the answer blank. Numerical Answers Expected


1
Expert's answer
2020-07-08T18:02:56-0400

3.

Given:

no. of goldfish in a tank = 22 = mass

volume of the tank = 56 cubic feet

So

Density = Mass/Volume

= 22/56

= 0.39 goldfish per cubic foot.

Answer: 0.39 goldfish per cubic foot.


4.

Given:

Density of Tinyville = 16.2 Tinys per acre.

Population in Tinyville = 405 Tinys.

Let x be total acres of Tinyville.

So

Tinys living in 1 acre = 16.2,

total Tinys living in Tinyville = 405.

So, x = 405/16.2 = 25 acres.

Answer: 25 acres.